OpenStudy (anonymous):

Help me find the inverse function!!! It is 2 am and I need to sleep!!!

OpenStudy (anonymous):

\[f(x) = \left| x \right| - 4\]

jimthompson5910 (jim_thompson5910):

Since |x| is NOT one-to-one, it does NOT have an inverse function. However, you can restrict the domain to create an inverse do they provide the restriction?

OpenStudy (anonymous):

Can you write the equation even though it's not a function?

jimthompson5910 (jim_thompson5910):

do they say something like "x > 0" or "x < 0" ?

OpenStudy (zzr0ck3r):

The inverse relation would be given by \(y=|x|+4\iff y-4=|x| \iff x=4-y \) or \(x=y-4\)

OpenStudy (zzr0ck3r):

But as Jim said, its not a function